# ShadowKV: KV Cache in Shadows for High-Throughput Long-Context LLM Inference

### v1

@@ Abstract
With the widespread deployment of long-context large language models (LLMs), there has been a growing demand for efficient support of high-throughput inference. However, as the key-value (KV) cache expands with the sequence length, the increasing memory footprint and the need to access it for each token generation both result in low throughput when serving long-context LLMs. While various dynamic sparse attention methods have been proposed to speed up inference while maintaining generation quality, they either fail to sufficiently reduce GPU memory consumption or introduce significant decoding latency by offloading the KV cache to the CPU. We present ShadowKV, a high-throughput long-context LLM inference system that stores the low-rank key cache and offloads the value cache to reduce the memory footprint for larger batch sizes and longer sequences. To minimize decoding latency, ShadowKV employs an accurate KV selection strategy that reconstructs minimal sparse KV pairs on-the-fly. By evaluating ShadowKV on a broad range of benchmarks, including RULER, LongBench, and Needle In A Haystack, and models like Llama-3.1-8B, Llama-3-8B-1M, GLM-4-9B-1M, Yi-9B-200K, Phi-3-Mini-128K, and Qwen2-7B-128K, we demonstrate that it can support up to 6 larger batch sizes and boost throughput by up to 3.04 on an A100 GPU without sacrificing accuracy, even surpassing the performance achievable with infinite batch size under the assumption of infinite GPU memory. The code is available at https://github.com/bytedance/ShadowKV.
@@ 1 Introduction
Large language models (LLMs) have increasingly demonstrated their ability to scale and handle long contexts [34, 30, 2, 49], enabling them to tackle complex tasks like multi-document question answering and information retrieval from extensive contexts of up to 1M tokens [2, 53]. However, efficiently serving these long-context LLMs presents challenges related to the key-value (KV) cache [31, 13], which stores previous key-value activations to avoid re-computation. As the KV cache scales with sequence length, its growing memory footprint and the need to access it for each token generation lead to low throughput during long-context LLM inference. To address these issues, KV cache eviction or sparse attention methods have been widely explored.
@@
However, existing methods face three primary limitations: accuracy degradation, inadequate memory reduction, and significant decoding latency overhead. KV cache eviction strategies [67, 66] aim to reduce the memory footprint by discarding KV pairs based on specific policies, but they often result in information loss and accuracy degradation in tasks such as multi-turn conversations [60, 47]. Dynamic sparse attention methods [48] preserve all KV pairs on the GPU and accelerate inference by computing attention with selected KV pairs. However, this line of work does not mitigate the memory footprint, thereby limiting the batch size and preventing accommodation of extremely long contexts (e.g., 1M tokens). A naive solution based on sparse attention involves offloading the KV cache to the CPU to reduce memory usage [26, 16]. Nonetheless, this approach incurs significant overhead due to the latency of fetching the selected sparse KV pairs from the CPU during decoding.
@@
Consequently, an ideal effective system for long-context LLM inference with sparse attention should: (i) reduce GPU memory usage, (ii) minimize inference latency, and (iii) maintain accuracy within limited sparse KV cache budgets. Fortunately, we can potentially overcome these challenges by leveraging our discovery that pre-Rotary Position Embedding [45] (RoPE) keys are exceptionally low-rank compared to the layer inputs, post-RoPE keys, values, key weight matrix, and value weight matrix, as indicated in Figure 1(a). Furthermore, our analysis in Figure 1(b) reveals that pre-RoPE keys lack significant similarities in low-rank subspaces across different sequences, while a sequence and its continuation tend to strongly share low-rank subspaces, enabling high compression rates within each sequence. Motivated by these findings, we have developed two key insights that pave the way for the design of an applicable system, detailed in Section 3.
@@
Low-rank Keys and Offloaded Values for Storage: In long-context LLM inference, the quadratic scaling of attention computation with sequence length makes the linear cost of low-rank decomposition during pre-filling negligible, as illustrated in Figure 1(c)111In practical scenarios, the key cache can be offloaded to the CPU to perform SVD asynchronously or precomputed and stored as part of the prefix cache [23].. To reduce memory footprint, we retain the low-rank pre-RoPE key cache on the GPU and offload the value cache to the CPU since the value cache does not exhibit low-rank properties, minimizing memory footprint without sacrificing accuracy. During decoding with sparse attention, we employ CUDA multi-streams to overlap the recovery of the selected key cache with the fetching of the corresponding value cache. This approach conceals key cache reconstruction and reduces data fetching overhead by compared to the naive offloading strategy, thereby decreasing the latency of sparse attention during decoding.
@@
Accurate KV Selection for Fast Decoding: To further reduce decoding latency in sparse attention, we propose an accurate KV selection method that maintains accuracy with minimal sparse budgets (1.56%). Our analysis reveals that most post-RoPE keys exhibit high cosine similarity with adjacent tokens, enabling chunk-level approximations for selecting important tokens. A minimal number of outlier chunks (0.3%), which are more challenging to approximate (Figure 3(b)), are stored as static cache on the GPU to preserve accuracy. As shown in Figure 2, our method outperforms the naive sparse attention approach [48] and achieves higher sparsity, accelerating decoding.
@@
In Section 5.2, we demonstrate that ShadowKV can support 6 larger batch sizes and boost throughput by 3.04 compared to small batches on an A100 using Llama-3.1-8B, with each sample having a context length of 122K. We also present results across different models and context lengths, increasing throughput up to 2.97 for Llama-3-8B-1M, 2.56 for GLM-4-9B-1M, and 2.66 for Yi-9B-200K, even surpassing infinite batch size under the assumption of infinite GPU memory.
@@ 2 Related Works
To reduce memory footprint, eviction-based strategies keep a fixed size of KV cache to store the critical token KV pairs and discard unnecessary tokens. StreamingLLM [57] addresses the limitations of window attention by retaining attention sinks and recent KV pairs. H2O [67] introduces a low-cost eviction policy, updating the KV cache based on cumulative attention scores. LESS [11] accumulates evicted token information by a constant-sized low-rank cache, which allows partial access to previously evicted information, along with tokens maintained by a sparse policy. SnapKV [29] uses the local window of prompts to select important tokens for future generations. However, they suffer from performance degradation and information loss since the evicted tokens will never be recovered.
@@
This line of work retains all KV cache but performs dynamic sparse attention within selected KV pairs to reduce inference latency. SparQ [41] uses the norm of the query to decide an important subset of the key cache’s channels to calculate a metric to select relevant tokens. Quest [48] segments tokens into pages and selects pages by approximating the highest attention within each page. Loki [43] performs principal component analysis on the key cache using a calibration dataset, selecting tokens based on attention scores computed in low-dimensional space. TriForce [46] combines sparse attention with speculative decoding [27] for lossless acceleration. InfiniGen [26] offloads all KV cache to the CPU and prefetches essential entries during decoding.
@@ 3.1 Low-Rank Keys and Offloaded Values for Storage
To reduce memory footprint, the low-rank nature of the KV cache has been explored by recent studies [9, 58, 5]. However, these methods focus on data-independent decomposition, either requiring training or achieving limited compression rates.
@@
In our study, by conducting SVD on the model weights , , the input , the pre-/post-RoPE key cache, and the value cache of Llama-3.1-8B, we visualize the relative singular value distributions in Figure 1(a) together with the accuracy in Figure 3(a). As we observed, pre-RoPE keys have the lowest rank and can be compressed by 6 without performance degradation.
We also identify striking dynamic and static behaviors in low-rank keys between and within sequences, inspired by a related investigation in FFN layers [10]. Analogous to cosine similarity, we define to be the similarity metric between low-rank subspaces of two rank- projection matrices, and , where is the Frobenius inner product222Since and are projection matrices, their squared Frobenius norms are the sum of their singular values which consist of 1’s and 0’s, i.e., . Thus, by Cauchy-Schwarz, . Additionally, by the cyclic property of trace and positive semidefiniteness of projection matrices. Together, this shows , maximized or minimized when the projection matrices project onto identical or orthogonal subspaces, respectively.. In our case with truncated SVDs of pre-RoPE keys, let have rank- truncated SVDs, and , respectively, where , and similarly for , , and . Then, can measure the similarity between the low-rank subspaces of the two right singular matrices. Depicted in Figure 1(b), pre-RoPE keys between sequences do not strongly share similar low-rank subspaces, but extensions of the same sequence do.
@@ 3.2 Accurate KV Selection for Fast Decoding
We found most post-RoPE key cache exhibits spatial locality, with high cosine similarity to adjacent tokens, except for a few outliers. To quantify this, we conducted inference experiments on 128K contexts. We divided the post-RoPE keys into chunks of eight tokens and visualized the minimum cosine similarity between the chunk’s mean and its key cache, as shown in Figure 3(b). The results indicate that, apart from a few outliers, there is generally high cosine similarity, suggesting the mean values can serve as landmarks to approximate attention well within normal chunks.
@@
This finding suggests that for the majority of chunks, we can maintain the mean value as compressed landmarks to select minimal important KV pairs (1.56%) accurately during decoding. Outlier chunks, which may contain dense or critical information and are difficult to approximate, are retained to ensure accuracy. Given their relatively small number (0.2–0.3%), storing them on the GPU is feasible without affecting memory capacity. Furthermore, as shown in Figure 3(c), considering the temporal locality of the KV cache, a cache policy [64] can be leveraged to further reduce the latency overhead by 60% during decoding with optimized kernels [35].
@@ 4 ShadowKV
In this section, we introduce ShadowKV, a high-throughput long-context LLM inference system. We first elaborate our algorithm in Section 4.1, covering both the pre-filling and decoding phases. Subsequently, in Section 4.2, we discuss the concept of theoretical equivalent bandwidth to illustrate the benefits of our approach.
@@ 4.1 Algorithm
During the pre-filling phase, we optimize GPU memory usage by performing low-rank compression on the key cache of each layer and offloading values to the CPU. Specifically, as demonstrated in Algorithm 1 and Figure 4, we apply SVD on the pre-RoPE key cache and store only the low-rank representations for each layer. Post-RoPE key cache is segmented into chunks, with the mean of each chunk computed as landmarks. By computing the cosine similarity within these chunks, we identify poorly approximated tokens as outliers. This small set of outliers is gathered and stored on the GPU as the static cache, while the remaining key cache is maintained as compact landmarks, with the corresponding values offloaded to the CPU memory.
@@
For incoming queries, we first compute the approximate attention scores using the landmarks. As detailed in Algorithm 2, by identifying the top-k scoring chunk indices, the corresponding values are retrieved from the CPU, and the key cache is simultaneously reconstructed from low-rank projections, effectively concealing the construction of the key cache. Based on the insight that the KV cache has temporal locality, we build cache-aware CUDA kernels, reducing computation and value fetching by 60%. As shown in Figure 5, we conduct an index scan to detect the missed chunks and only rebuild the necessary KV pairs on-the-fly.
@@
Based on our observations in Section 3.1, future pre-RoPE keys within a sequence reside in a shared low-rank subspace with the context. As a result, an extension of our algorithm would be to store generated tokens as low-rank states using the same projections obtained from pre-filling to reduce the memory usage for future generations333If is the right singular matrix calculated from the SVD of pre-RoPE context keys , new pre-RoPE keys can be stored as and projected back up with when needed.. For simplicity, we exclude it from the implementation.

### v2

@@ Abstract
With the widespread deployment of long-context large language models (LLMs), there has been a growing demand for efficient support of high-throughput inference. However, as the key-value (KV) cache expands with the sequence length, the increasing memory footprint and the need to access it for each token generation both result in low throughput when serving long-context LLMs. While various dynamic sparse attention methods have been proposed to speed up inference while maintaining generation quality, they either fail to sufficiently reduce GPU memory consumption or introduce significant decoding latency by offloading the KV cache to the CPU. We present ShadowKV, a high-throughput long-context LLM inference system that stores the low-rank key cache and offloads the value cache to reduce the memory footprint for larger batch sizes and longer sequences. To minimize decoding latency, ShadowKV employs an accurate KV selection strategy that reconstructs minimal sparse KV pairs on-the-fly. By evaluating ShadowKV on a broad range of benchmarks, including RULER, LongBench, and Needle In A Haystack, and models like Llama-3.1-8B, Llama-3-8B-1M, GLM-4-9B-1M, Yi-9B-200K, Phi-3-Mini-128K, and Qwen2-7B-128K, we demonstrate that it can support up to 6 larger batch sizes and boost throughput by up to 3.04 on an A100 GPU without sacrificing accuracy, even surpassing the performance achievable with infinite batch size under the assumption of infinite GPU memory.
[…]
@@ 1 Introduction
Large language models (LLMs) have increasingly demonstrated their ability to scale and handle long contexts [2, 46, 32, 28], enabling them to tackle complex tasks like multi-document question answering and information retrieval from extensive contexts of up to 1M tokens [2, 49]. However, efficiently serving these long-context LLMs presents challenges related to the key-value (KV) cache [11, 29], which stores previous key-value activations to avoid re-computation. As KV cache scales with sequence length, its growing memory footprint and the need to access it for each token generation lead to low throughput during long-context LLM inference. To address these, KV eviction or sparse attention methods have been widely explored.
@@
However, existing methods face three primary limitations: accuracy degradation, inadequate memory reduction, and significant decoding latency overhead. KV cache eviction strategies [63, 62] aim to reduce the memory footprint by discarding KV pairs based on specific policies, but they often result in information loss and accuracy degradation in tasks such as multi-turn conversations [55, 44]. Dynamic sparse attention methods [45] preserve all KV pairs on the GPU and accelerate inference by computing attention with selected KV pairs. However, this line of work does not mitigate the memory footprint, thereby limiting the batch size and preventing accommodation of extremely long contexts (e.g., 1M tokens). A naive solution based on sparse attention involves offloading the KV cache to the CPU to reduce memory usage [22, 14]. Nonetheless, as shown in Figure 4, this approach incurs significant overhead due to the latency of fetching the selected sparse KV pairs from the CPU during decoding.
@@
Consequently, an ideal system for long-context LLM inference with sparse attention should: (i) reduce GPU memory usage, (ii) minimize inference latency, and (iii) maintain accuracy within limited sparse KV cache budgets. Fortunately, we can potentially overcome these challenges by leveraging our discovery that pre-Rotary Position Embedding [41] (RoPE) keys are exceptionally low-rank compared to the layer inputs, post-RoPE keys, values, key weight matrix, and value weight matrix, as indicated in Figure 1. Unlike prior works that apply low-rank approximations to weights [5, 23], we directly compress pre-RoPE key cache, achieving higher accuracy. Our analysis reveals that pre-RoPE keys lack significant similarities in low-rank subspaces across different sequences, while a sequence and its continuation tend to strongly share low-rank subspaces, enabling high compression rates within each sequence. Motivated by these findings, we developed two key insights that pave the way for the design of an applicable system, detailed in Section 3.
@@
Low-rank Keys and Offloaded Values for Storage: In long-context LLM inference, the quadratic scaling of attention computation with sequence length makes the linear cost of low-rank decomposition during pre-filling negligible, as illustrated in Figure 1111In practical scenarios, the key cache can be offloaded to the CPU to perform SVD asynchronously or precomputed and stored as part of the prefix cache [19].. To reduce memory footprint, we retain the low-rank pre-RoPE key cache on the GPU and offload the value cache to the CPU since the value cache does not exhibit low-rank properties, minimizing memory footprint without sacrificing accuracy. During decoding with sparse attention, we employ CUDA multi-streams to overlap the recovery of the selected key cache with the fetching of the corresponding value cache. This approach conceals key cache reconstruction and reduces data fetching overhead by compared to the naive offloading strategy, thereby decreasing the decoding latency of sparse attention.
@@
Accurate KV Selection for Fast Decoding: To further reduce decoding latency in sparse attention, we propose an accurate KV selection method that maintains accuracy with minimal number of selected tokens (i.e., the K of TopK), which we refer to as sparse budgets (1.56%). Our analysis reveals that most post-RoPE keys exhibit high cosine similarity with adjacent tokens, enabling chunk-level approximations for selecting important tokens. A minimal number of outlier chunks (0.3%), which are more challenging to approximate (Figure 5), are stored as static cache on the GPU to preserve accuracy. As shown in Figure 2, our method outperforms the naive sparse attention approach [45] and achieves higher sparsity, accelerating decoding.
@@
In Section 5.2, we demonstrate that ShadowKV can support 6 larger batch sizes and boost throughput by up to 3.04 compared to small batches on an A100 using Llama-3.1-8B. We also present results across different models and context lengths, increasing throughput up to 2.97 for Llama-3-8B-1M, 2.56 for GLM-4-9B-1M, and 2.66 for Yi-9B-200K, even surpassing infinite batch size under the assumption of infinite GPU memory.
@@ 2 Related Works
To reduce memory footprint, eviction-based strategies keep a fixed size of KV cache to store the critical token KV pairs and discard unnecessary tokens. StreamingLLM [52] addresses the limitations of window attention by retaining attention sinks and recent KV pairs. H2O [63] introduces a low-cost eviction policy, updating the KV cache based on cumulative attention scores. LESS [10] accumulates evicted token information by a constant-sized low-rank cache, which allows partial access to evicted information, along with tokens maintained by a sparse policy. SnapKV [26] uses the local window of prompts to select important tokens for future generations. However, they suffer from performance degradation and information loss since the evicted tokens will never be recovered.
@@
This line of work retains all KV cache but performs dynamic sparse attention within selected KV pairs to reduce inference latency. SparQ [36] uses the norm of the query to decide an important subset of the key cache’s channels to calculate a metric to select relevant tokens. Quest [45] segments tokens into pages and selects pages by approximating the highest attention within each page. Loki [39] performs principal component analysis on key caches using a calibration dataset, selecting tokens based on attention scores computed in low-dimensional space. TriForce [42] combines sparse attention with speculative decoding [24] for lossless acceleration. InfiniGen [22] offloads the entire KV cache to the CPU and prefetches essential entries using predefined projections via SVD for KV selection. In contrast, ShadowKV employs an online, prompt-dependent SVD for key cache compression, reducing data fetching. As shown in Figure 4, ShadowKV scales better, as overlapping KV fetching and computation becomes challenging with larger KV caches. Additionally, InfiniGen suffers performance drops due to inaccurate prefetching KV selection.
@@ 3.1 Low-Rank Keys and Offloaded Values for Storage
To reduce memory footprint, the low-rank nature of the KV cache has been explored by recent studies [8, 53, 5, 37, 57, 27]. However, these methods focus on data-independent decomposition, either requiring training or achieving limited compression rates.
@@
In our study, we visualize the relative singular value distributions of the model weights , , the input , the pre-/post-RoPE key cache, and the value cache of Llama-3.1-8B by conducting SVD, as shown in Figure 1. In Figure 5, we further analyze the accuracy impact. We observe that pre-RoPE keys exhibit the lowest rank, allowing for compression without performance degradation.
We also identify striking dynamic and static behaviors in low-rank keys between and within sequences, inspired by a related investigation in FFN layers [9]. Analogous to cosine similarity, we define to be the similarity metric between low-rank subspaces of two rank- projection matrices, and , where is the Frobenius inner product222Since and are projection matrices, their squared Frobenius norms are the sum of their singular values which consist of 1’s and 0’s, i.e., . Thus, by Cauchy-Schwarz, . Additionally, by the cyclic property of trace and positive semidefiniteness of projection matrices. Together, this shows , maximized or minimized when the projection matrices project onto identical or orthogonal subspaces, respectively.. In our case with truncated SVDs of pre-RoPE keys, let have rank- truncated SVDs, and , respectively, where , and similarly for , , and . Then, can measure the similarity between the low-rank subspaces of the two right singular matrices. Depicted in Figure 1, pre-RoPE keys between sequences do not strongly share similar low-rank subspaces, but extensions of the same sequence do. Thus, applying low-rank approximations to weights alone degrades performance.
@@ 3.2 Accurate KV Selection for Fast Decoding
We found most post-RoPE key cache exhibits spatial locality, with high cosine similarity to adjacent tokens, except for a few outliers. To quantify this, we conducted inference experiments on 128K contexts. We divided the post-RoPE keys into chunks of eight tokens and visualized the minimum cosine similarity between the chunk’s mean and its key cache. The results indicate that, apart from a few outliers, there is generally high cosine similarity, suggesting the mean values can serve as landmarks to approximate attention well within normal chunks.
@@
This finding suggests that for the majority of chunks, we can maintain the mean value as compressed landmarks to select minimal important KV pairs (1.56%) accurately during decoding. Outlier chunks, which may contain dense or critical information and are difficult to approximate, are retained to ensure accuracy. Given their relatively small number (0.2–0.3%), storing them on the GPU is feasible without affecting memory capacity. Furthermore, as shown in Figure 5, considering the temporal locality of the KV cache—meaning that the KV pairs selected by the queries of two adjacent decoding steps have a high repetition rate, a cache policy [60] can be leveraged to further reduce the latency overhead by 60% during decoding with optimized CUDA kernels.
@@ 4 ShadowKV
In this section, we introduce ShadowKV, a novel high-throughput long-context LLM inference system. We first elaborate our algorithm in Section 4.1, covering both the pre-filling and decoding phases. Subsequently, in Section 4.2, we discuss the concept of theoretical equivalent bandwidth to illustrate the benefits of our approach.
@@ 4.1 Algorithm
During the pre-filling phase, we optimize GPU memory usage by performing low-rank compression on the key cache of each layer and offloading values to the CPU. Specifically, as demonstrated in Algorithm 1, we apply SVD on the pre-RoPE key cache and store only the low-rank representations for each layer. Post-RoPE key cache is segmented into chunks, with the mean of each chunk computed as landmarks. By computing the cosine similarity within these chunks, we identify poorly approximated tokens as outliers. This small set of outliers is gathered and stored on the GPU as the static cache, while the remaining key cache is maintained as compact landmarks, with the corresponding values offloaded to the CPU memory.
@@
For incoming queries, we first compute the approximate attention scores using the landmarks. As detailed in Algorithm 2, by identifying the top-k scoring chunk indices, the corresponding values are retrieved from the CPU, and the key cache is simultaneously reconstructed from low-rank projections, effectively concealing the construction of the key cache. Based on the insight that the KV cache has temporal locality, we conduct an index scan to detect the missed chunks and only rebuild the necessary KV pairs on-the-fly, reducing computation and data fetching by 60% with optimized CUDA kernels.
@@
Based on our observations in Section 3.1, future pre-RoPE keys within a sequence reside in a shared low-rank subspace with the context. As a result, an extension of our algorithm would be to store generated tokens as low-rank states using the same low-rank projections obtained from pre-filling to reduce the memory usage for the future generations333If is the right singular matrix calculated from the SVD of pre-RoPE context keys , new pre-RoPE keys can be stored as and projected back up with when needed.. We evaluate it and include the results in Section 7.1.
Methods | S1 | S2 | MK1 | MK2 | MQ | MV | QA-1 | QA-2 | VT | FWE | Avg. | NQA | MQA | HQA | MQue | DRead | GRep | SAM | PRetr | LCC | Avg. |
Llama-3-8B-1M | 100.00 | 100.00 | 98.96 | 98.96 | 98.96 | 95.57 | 75.00 | 48.96 | 78.54 | 71.85 | 86.68 | 18.98 | 41.84 | 36.79 | 21.47 | 31.93 | 34.18 | 35.96 | 81.50 | 56.07 | 39.86 |
Loki | 18.75 | 1.04 | 2.08 | 0.00 | 1.56 | 0.78 | 4.17 | 13.54 | 26.04 | 25.35 | 9.33 | 2.26 | 10.19 | 5.48 | 3.16 | 12.17 | 28.97 | 7.84 | 40.52 | 31.44 | 15.78 |
Loki (V) | 41.67 | 6.25 | 37.50 | 1.04 | 8.07 | 30.73 | 10.42 | 19.79 | 51.67 | 37.50 | 24.46 | 3.20 | 21.01 | 12.41 | 3.86 | 17.07 | 31.24 | 16.23 | 52.57 | 38.10 | 21.74 |
InfiniGen | 100.00 | 98.96 | 84.38 | 53.13 | 63.28 | 54.95 | 65.63 | 48.96 | 81.67 | 50.35 | 70.13 | 14.39 | 31.46 | 33.63 | 17.94 | 26.65 | 27.38 | 21.97 | 74.30 | 38.58 | 31.81 |
InfiniGen (V) | 100.00 | 98.96 | 96.88 | 76.04 | 81.25 | 77.08 | 67.71 | 50.00 | 81.67 | 53.47 | 78.31 | 17.83 | 36.08 | 35.28 | 19.64 | 28.39 | 29.28 | 28.12 | 74.85 | 45.53 | 35.00 |
Quest | 100.00 | 100.00 | 98.96 | 77.08 | 97.65 | 93.49 | 60.42 | 50.00 | 77.08 | 65.63 | 82.03 | 20.13 | 36.63 | 35.00 | 18.14 | 24.55 | 27.11 | 35.63 | 79.00 | 53.64 | 36.65 |
Quest (V) | 100.00 | 100.00 | 98.96 | 85.42 | 97.92 | 95.49 | 70.83 | 46.88 | 78.75 | 65.63 | 83.99 | 17.26 | 39.51 | 36.78 | 18.71 | 26.41 | 29.49 | 35.80 | 79.50 | 60.05 | 38.17 |
ShadowKV | 100.00 | 100.00 | 97.92 | 98.96 | 96.88 | 95.83 | 72.92 | 52.08 | 81.67 | 72.57 | 86.88 | 17.17 | 39.73 | 38.29 | 21.08 | 31.77 | 31.62 | 35.87 | 80.00 | 63.93 | 39.94 |
GLM-4-9B-1M | 100.00 | 100.00 | 94.79 | 87.50 | 99.74 | 93.75 | 67.71 | 55.21 | 97.29 | 72.22 | 86.82 | 25.44 | 51.09 | 58.67 | 39.61 | 32.04 | 29.97 | 40.31 | 99.00 | 58.02 | 48.24 |
Loki | 71.88 | 27.08 | 22.92 | 2.08 | 9.90 | 11.46 | 28.13 | 27.08 | 31.04 | 54.17 | 28.57 | 5.82 | 30.60 | 22.73 | 9.20 | 30.09 | 30.35 | 22.70 | 98.92 | 40.77 | 32.35 |
Loki (V) | 96.88 | 55.21 | 56.25 | 18.75 | 51.04 | 50.52 | 45.83 | 39.58 | 72.71 | 59.72 | 54.65 | 10.89 | 44.97 | 45.44 | 23.51 | 32.07 | 30.56 | 35.34 | 99.50 | 50.27 | 41.39 |
InfiniGen | 100.00 | 93.75 | 82.29 | 0.00 | 79.43 | 60.16 | 57.29 | 53.13 | 92.71 | 57.29 | 67.60 | 23.67 | 46.31 | 55.69 | 33.91 | 27.49 | 25.44 | 33.48 | 91.83 | 36.96 | 41.64 |
InfiniGen (V) | 100.00 | 96.88 | 87.50 | 7.29 | 95.31 | 75.26 | 56.25 | 54.17 | 95.63 | 60.76 | 72.91 | 25.63 | 48.44 | 57.23 | 36.94 | 29.77 | 26.67 | 36.64 | 93.58 | 46.69 | 44.62 |
Quest | 100.00 | 95.83 | 90.62 | 54.17 | 94.01 | 76.30 | 55.21 | 52.08 | 95.83 | 64.58 | 77.86 | 23.81 | 44.53 | 56.41 | 35.49 | 23.54 | 21.73 | 37.39 | 87.00 | 43.80 | 41.52 |
Quest (V) | 100.00 | 96.88 | 93.75 | 72.92 | 95.83 | 83.07 | 56.25 | 53.13 | 96.88 | 65.97 | 81.47 | 26.00 | 46.32 | 57.54 | 36.42 | 24.58 | 24.52 | 37.71 | 93.50 | 46.52 | 43.68 |
ShadowKV | 100.00 | 100.00 | 95.83 | 83.33 | 98.70 | 87.76 | 69.79 | 55.21 | 97.50 | 68.06 | 85.62 | 26.50 | 51.31 | 59.09 | 38.87 | 32.92 | 28.54 | 38.70 | 96.50 | 58.55 | 47.89 |
Llama-3.1-8B | 100.00 | 100.00 | 98.96 | 91.67 | 98.96 | 95.31 | 82.29 | 47.92 | 68.96 | 71.18 | 85.53 | 31.56 | 55.10 | 57.65 | 29.46 | 35.26 | 34.45 | 29.84 | 100.00 | 67.31 | 48.96 |
Loki | 68.75 | 32.29 | 32.29 | 20.83 | 42.71 | 28.65 | 41.67 | 33.33 | 24.79 | 29.86 | 35.52 | 2.31 | 18.89 | 10.64 | 5.47 | 19.30 | 31.16 | 15.91 | 94.88 | 44.60 | 27.02 |
Loki (V) | 95.83 | 36.46 | 57.29 | 62.50 | 77.86 | 70.83 | 69.79 | 39.58 | 35.21 | 37.50 | 58.29 | 3.93 | 38.59 | 22.85 | 12.96 | 27.43 | 32.22 | 26.43 | 98.25 | 56.11 | 35.42 |
InfiniGen | 100.00 | 77.08 | 78.13 | 13.54 | 58.07 | 47.40 | 65.63 | 41.67 | 60.83 | 50.35 | 59.27 | 27.23 | 52.72 | 53.89 | 26.81 | 27.72 | 29.61 | 24.42 | 98.93 | 56.89 | 44.25 |
InfiniGen (V) | 100.00 | 88.54 | 87.50 | 26.04 | 79.43 | 77.08 | 72.92 | 43.75 | 57.08 | 55.21 | 68.76 | 29.73 | 53.47 | 55.11 | 28.72 | 28.55 | 31.42 | 26.76 | 99.17 | 62.66 | 46.18 |
Quest | 100.00 | 98.96 | 97.92 | 34.38 | 93.49 | 88.54 | 70.83 | 44.79 | 65.63 | 68.40 | 76.29 | 29.70 | 49.04 | 53.96 | 27.18 | 27.16 | 30.43 | 29.85 | 98.50 | 57.35 | 44.80 |
Quest (V) | 100.00 | 98.96 | 98.96 | 56.25 | 95.83 | 90.63 | 76.04 | 46.88 | 66.25 | 67.36 | 79.72 | 30.02 | 53.97 | 56.39 | 27.06 | 29.06 | 31.65 | 30.23 | 99.00 | 63.89 | 46.81 |
ShadowKV | 100.00 | 100.00 | 100.00 | 83.33 | 97.92 | 92.19 | 81.25 | 48.96 | 67.08 | 64.93 | 83.57 | 30.93 | 55.20 | 57.32 | 29.13 | 31.85 | 32.79 | 30.40 | 99.50 | 66.03 | 48.13 |